# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

### v3

@@ 5.2 Flexible Rollout Strategies
As demonstrated in Figure 2(b) and Figure 5, VICON can select appropriate timestep strides based on each dataset’s characteristics. While PDEBench-Comp-LowVis and PDEBench-Comp-HighVis perform best with single-step rollout, PDEArena-Incomp achieves optimal results with a stride of 3 using our flexible-step rollout strategy.
This dataset-dependent performance is expected: PDEBench-Comp-LowVis and PDEBench-Comp-HighVis record data with a larger timestep size and fewer total frames, making multi-stride predictions challenging. Conversely, PDEArena-Incomp records data with smaller timesteps and benefits from multi-stride prediction, as this approach reduces the rollout steps and error accumulation. As a result, VICON achieves an 11.4% error reduction in PDEArena-Incomp by selecting a balanced stride ($s_{\text{max}}=3$).
@@ 5.3 Robustness to Imperfect Temporal Measurements
Real-world experimental measurements frequently suffer from imperfections—sampling rates may differ from training set, and frames can be missing due to device errors. Sequence-to-sequence models like MPP and DPOT require either retraining with new data or interpolation that introduces noise. In contrast, VICON elegantly handles such imperfections by forming context pairs from available frames only (Figure 1(c)). The algorithm for generating pairs with imperfect measurements appears in Section C.2.
We evaluate VICON against baselines on all benchmarks under two scenarios: (1) half sampling rate (dropping every other frame) and (2) random frame dropping (removing 1-3 frames per sequence). Results for the PDEBench-Comp-LowVis dataset in Figure 2(c) (with complete results in Figure 8) show that VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% in baselines—demonstrating remarkable robustness to measurement imperfections.
@@ 5.4 Turbulence Kinetic Energy Analysis
Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of [49]), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON, DPOT, and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming MPP’s error of $0.049$, while achieving performance comparable to DPOT’s error of $0.012$. Visualizations of TKE errors are presented in Figure 6 in Appendix D.
@@ 5.5 Benefits of Multi-Physics Joint Training
We empirically examine whether VICON benefits from a multi-physics training by comparing two strategies: (1) training a single model jointly on all three datasets versus (2) training separate models specialized for individual datasets. For fair comparison, we maintained identical batch sizes across all trainings while adjusting the steps for separate training to be slightly more than one-third of the joint training duration, ensuring comparable total computational costs between approaches.
As shown in Figure 7, joint training significantly outperforms separate training across all three datasets. We attribute this performance gain to the underlying physical similarities across these fluid dynamics problems, where exposure to diverse yet related flow patterns enhances the model’s generalization ability.
@@ 5.6 Computational Efficiency
In Figure 2(d) and Table 11, we compare the computational resources required by VICON and baselines. Our model demonstrates superior efficiency across multiple metrics. Compared to MPP, VICON requires approximately one-third of the inference time per frame while using 75% of the total parameters. VICON also outperforms DPOT, requiring 28% fewer parameters and 28% less inference time.
6 Conclusion and Future Works
We present VICON, a vision-transformer-based in-context operator network that efficiently processes dense physical fields through patch-wise operations. VICON overcomes the computational burden of original in-context operator networks in higher spatial dimensions while preserving the flexibility to extract multiphysics dynamics from few-shot contexts.
@@ 6 Conclusion and Future Works
Our comprehensive experiments demonstrate that VICON achieves superior performance in long-term predictions, reducing the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times. The model supports flexible rollout strategies with varying timestep strides, enabling natural application to imperfect real-world measurements where sampling frequencies differ or frames are randomly dropped—scenarios where VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% degradation in baseline models.
Despite these advances, several challenges remain for future investigation. We emperically showed in Section 5.5 the benefits of multi-physics training, yet scaling to larger and more diverse datasets presents practical challenges. Specifically, generating high-fidelity physics simulation data across multiple domains requires both substantial computational resources and specialized domain expertise. Furthermore, the current approach does not yet extend to 3D applications, as token sequence length would grow cubically, exceeding our computational budget. The channel-union approach is also limited when incorporating domains beyond fluid dynamics with fundamentally different state variables. Finally, adapting VICON to handle irregular domains such as graphs or meshes would broaden its applications to areas like solid mechanics and molecular dynamics. Addressing these challenges will advance the promising paradigm of in-context learning for a broader range of physical systems.
Acknowledgment
@@ A.1 PDEArena-Incomp Dataset
The space-time domain is $[0,32]^{2}\times[18,105]$ where $dt=1.5$ and the space resolution is $128\times 128$. A scalar particle field is being transported with the fluids. The velocity fields satisfy Dirichlet boundary condition, and the scalar field satisfies Neumann boundary condition. The forcing term $\bm{f}$ is randomly sampled. The quantities of interest are the velocities and the scalar particle field.
@@ B.1 VICON Model Details
Here we provide key architectural parameters of VICON model implementation in Table 2.
@@ B.2 Training Details
We implement our method in PyTorch [50] and utilize data parallel training [51] across two NVIDIA RTX 4090 GPUs. We employ the AdamW optimizer with a cosine learning rate schedule that includes a linear warmup phase. We apply gradient clipping with a maximum norm of 1.0 to ensure training stability. All optimization parameters are detailed in Table 3.
@@ 2 S $\leftarrow$ [ ]
We demonstrate two rollout strategies in Tables 4 and 5, showing single-step and flexible-step strategies, respectively. In both cases, the initial frames span from time step 0 to 9, and we aim to predict the trajectory up to time step 20.
@@
As shown in Table 5, the flexible-step strategy initially uses smaller strides to build sufficient examples, then employs maximum strides for later rollouts, as detailed in Section 4.5. We note that repeated in-context examples appear in Table 5, which is common when the maximum stride is large and the initial frames cannot form enough examples. While the number of in-context examples in VICON is flexible and our model can accommodate fewer examples than the designed length, our preliminary experiments indicate that the model performs better with more in-context examples, even when some examples are repeated.
@@ 30       Pdt $\leftarrow$ P[s]
Tables 6 and 7 illustrate our adaptive strategies when frames 2, 5, and 9 are missing from the initial sequence (with all other settings identical to those in Section C.1). The tables demonstrate single-step and flexible-step (max stride: 3) approaches, respectively, highlighting how our algorithm dynamically selects appropriate example pairs to maintain prediction capability despite missing some temporal samplings.
@@ D.1 Ablation Studies
We conducted ablation studies on patch resolution by varying patch sizes (8, 16, 32, 64) to find a balance between spatial granularity and computational resource constraint. While smaller patches theoretically capture finer details, they generate longer token sequences, hitting memory caps due to transformer’s quadratic complexity. For patches smaller than $16\times 16$, we had to reduce token dimensions (1024→512) and feedforward dimensions (from 2048 to 1024), while patches below $8\times 8$ remained infeasible even with mixed precision training. Figure 4(a-c) shows that patch size 16 achieves optimal performance across all datasets—balancing sequence length and representational capacity. Performance degradation with coarser patches (32, 64) aligns with expectations, while degradation with finer patches stems from the necessary reduction in hidden dimensions, highlighting a fundamental challenge in scaling to 3D applications.
@@
We evaluated different combinations of positional encodings as shown in Figure 4(d-f). Our architecture employs two types of encodings: patch position encodings (for spatial relationships between patches) and function encodings (differentiating input/output in different pairs) (Section 4.2).
@@
We investigated how the number of in-context examples affects model performance, as shown in Figure 4(g-i). Following insights from the original ICON work [25], we evaluated context lengths of 6, 8, and 10 pairs, balancing performance against computational efficiency. The results show that 10 in-context examples deliver significantly better results for PDEBench-Comp-LowVis compared to shorter contexts, while for the remaining datasets, 10 and 8 in-context examples perform similarly (both outperforming 6 examples).
D.2 More Results and Visualizations
@@
Step 1 [1e-2] PDEArena-Incomp 11.01 11.01 5.97 6.17
@@
Step 5 [1e-2] PDEArena-Incomp 22.68 20.62 11.70 14.81
@@
Step 10 [1e-2] PDEArena-Incomp 35.67 30.53 20.74 27.89
@@
Last step [1e-2] PDEArena-Incomp 76.77 68.03 65.27 93.52
@@
All average [1e-2] PDEArena-Incomp 56.27 48.50 41.20 55.95
@@
Step 1 [1e-2] PDEArena-Incomp 5.63 5.63 3.02 3.12
@@
Step 5 [1e-2] PDEArena-Incomp 10.38 9.46 5.37 6.79
@@
Step 10 [1e-2] PDEArena-Incomp 14.65 12.55 8.55 11.49
@@
Last Step [1e-2] PDEArena-Incomp 16.50 14.56 14.03 19.47
@@
All average [1e-2] PDEArena-Incomp 16.26 14.31 11.86 15.77
@@
Tables 8 and 9 summarize the relative and absolute $L^{2}$ rollout errors across different timesteps for all evaluated models on the three datasets.
@@
Figure 5 illustrates VICON’s performance with varying timestep strides ($s_{\text{max}}=1,3,5,7$) across all three datasets, demonstrating the model’s ability to adapt to different temporal resolutions without retraining.
@@
Figure 6 displays the visualization of TKE fields for ground truth and model predictions, highlighting VICON’s superior ability to preserve critical turbulent flow structures compared to baseline approaches.
@@
Step 1 PDEArena-Incomp 11.10 9.66
@@
Step 5 PDEArena-Incomp 23.00 24.09
@@
Step 10 PDEArena-Incomp 36.18 41.66
@@
Last Step PDEArena-Incomp 77.81 123.0
@@
All average PDEArena-Incomp 56.26 76.35
@@
Figure 7 and Table 10 quantify the performance differences between jointly trained and separately trained models, demonstrating the significant advantages of multi-physics training across all datasets.
@@
Table 11 summarizes the computational resources required by each method, showing VICON’s advantages in terms of training cost, inference speed, and model parameter count compared to both DPOT and MPP baselines.
@@
Figure 8: Comparison of rollout errors (scale by std) with different input data noise levels across all datasets. Each column (left to right) shows VICON, DPOT, and MPP models, while each row (top to bottom) represents PDEBench-Comp-LowVis, PDEBench-Comp-HighVis, and PDEArena-Incomp datasets. For MPP and DPOT which require fixed $dt$ and context window, interpolation is used to generate missing frames, while VICON can directly handle irregular temporal data without interpolation.
@@
Figure 8 show the rollout results with imperfect temporal measurements.
@@
We compare the output of different models in Figure 9, Figure 10, and Figure 11. Figures 12 and 13 present additional visualizations of the VICON model outputs compared to ground truth and baseline predictions, highlighting the qualitative advantages of our approach.

### v4

@@ 5.2 Flexible Rollout Strategies
As demonstrated in fig.˜2(b) and fig.˜5, VICON can select appropriate timestep strides based on each dataset’s characteristics. While PDEBench-Comp-LowVis and PDEBench-Comp-HighVis perform best with single-step rollout, PDEArena-Incomp achieves optimal results with a stride of 3 using our flexible-step rollout strategy.
This dataset-dependent performance reflects the characteristics of each dataset and the challenges of learning larger timestep strides. PDEBench-Comp-LowVis and PDEBench-Comp-HighVis record data with larger timestep sizes, making the learning of larger stride predictions more challenging due to increased dynamics complexity. Additionally, the fewer number of total rollout steps (even with stride=1) in these datasets cancels out the benefits of reducing rollouts when using multi-stride. Conversely, PDEArena-Incomp records data with smaller timesteps, making both single- and multi-stride prediction learning more feasible. Under this condition, multi-stride reduces the total rollout steps and error accumulation, achieving an 11.4% error reduction with a balanced stride ($s_{\text{max}}=3$).
@@
5.3 Importance of In-Context Learning
To validate the importance of in-context learning, we compare VICON against a non-in-context ViT baseline (with similar architecture: 162M parameters, 12 layers, hidden dimension 1024 for attention, and 4096 for FFN, time stride $\Delta t=1$ for training) on the PDEArena-Incomp dataset. The results in table˜2 demonstrate that vanilla ViT cannot generalize to different timestep strides without in-context examples:
Table 2: Non-in-context ViT baseline comparison on PDEArena-Incomp dataset
Model Stride 1 Error Stride 2 Error
(1e-2) (1e-2)
ViT (Last step) 73.94 95.03
ViT (All average) 50.17 63.15
The significant performance degradation (95.03 vs 73.94 for last step error) demonstrates that vanilla ViT cannot generalize to different $\Delta t$ without in-context examples, validating our approach’s effectiveness.
Additionally, we evaluated the impact of incorrect context pairs on VICON’s performance using the PDEArena-Incomp dataset (table˜3).
Table 3: Impact of incorrect context pairs on VICON performance
Experiment Setting Average Rollout Error
Correct (stride 1) 0.55
Mix context pairs with stride 1 and 2 0.88
Context pairs with stride 2 0.92
Random noise context pairs 1.57
The clear performance degradation with incorrect context pairs confirms that VICON effectively utilizes in-context information rather than simply memorizing patterns.
@@ 5.4 Robustness to Imperfect Temporal Measurements
Real-world experimental measurements frequently suffer from imperfections—sampling rates may differ from training set, and frames can be missing due to device errors. Sequence-to-sequence models like MPP and DPOT require either retraining with new data or interpolation that introduces noise. In contrast, VICON elegantly handles such imperfections by forming context pairs from available frames only (fig.˜1(c)). The algorithm for generating pairs with imperfect measurements appears in section˜C.2.
We evaluate VICON against baselines on all benchmarks under two scenarios: (1) half sampling rate (dropping every other frame) and (2) random frame dropping (removing 1-3 frames per sequence). Results for the PDEBench-Comp-LowVis dataset in fig.˜2(c) (with complete results in fig.˜8) show that VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% in baselines—demonstrating remarkable robustness to measurement imperfections.
@@ 5.5 Turbulence Kinetic Energy Analysis
Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of [49]), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON, DPOT, and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming MPP’s error of $0.049$, while achieving performance comparable to DPOT’s error of $0.012$. Visualizations of TKE errors are presented in fig.˜6 in appendix˜D.
@@ 5.6 Benefits of Multi-Physics Joint Training
We empirically examine whether VICON benefits from multi-physics training by comparing two strategies: (1) training a single model jointly on all three datasets versus (2) training separate models specialized for individual datasets. For fair comparison, we maintained identical batch sizes across all trainings while adjusting the steps for separate training to be slightly more than one-third of the joint training duration, ensuring comparable total computational costs between approaches.
As shown in fig.˜7, joint training significantly outperforms separate training across all three datasets. We attribute this performance gain to the underlying physical similarities across these fluid dynamics problems, where exposure to diverse yet related flow patterns enhances the model’s generalization ability.
@@ 5.7 Computational Efficiency
In fig.˜2(d) and table˜14, we compare the computational resources required by VICON and baselines. Our model demonstrates superior efficiency across multiple metrics. Compared to MPP, VICON requires approximately one-third of the inference time per frame while using 75% of the total parameters. VICON also outperforms DPOT, requiring 28% fewer parameters and 28% less inference time.
It is important to note that, in theory, VICON incurs approximately 4× computational cost compared to non-paired, sequence-to-sequence baselines when all other settings (e.g., patch size) are identical, as it processes both condition and QoI, doubling the total token length with quadratic attention complexity. However, this computational overhead is justified by the significant performance improvements and unique capabilities (handling imperfect measurements, flexible timestep strides). Additionally, this cost can be reduced to 2× (linear) at inference time using KV caching, where all context pairs except the final pair remain fixed across rollout steps, allowing their key-value multiplications to be cached.
It is important to note that while post-training (such as fine-tuning) on a sequence-to-sequence model may achieve the same accuracy on one physical condition and save this 2× cost overhead, the fine-tuned model loses the ability to predict accurately when physical parameters may change on-the-fly during deployment, while VICON provides superior flexibility.
6 Conclusion and Future Work
@@
We present VICON, a vision-transformer-based in-context operator network that efficiently processes dense physical fields through patch-wise operations. VICON overcomes the computational burden of original in-context operator networks in higher spatial dimensions while preserving the flexibility to extract multi-physics dynamics from few-shot contexts. Our comprehensive experiments demonstrate that VICON achieves superior performance in long-term predictions, reducing the average last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times. The model supports flexible rollout strategies with varying timestep strides, enabling natural application to imperfect real-world measurements where sampling frequencies differ or frames are randomly dropped—scenarios where VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% degradation in baseline models.
Despite these advances, several challenges remain for future investigation. We empirically showed in section˜5.6 the benefits of multi-physics training, yet scaling to larger and more diverse datasets presents practical challenges. Specifically, generating high-fidelity physics simulation data across multiple domains requires substantial computational resources and specialized domain expertise including mesh generation, physical modeling, and numerical solver integration.
Furthermore, the current approach does not yet extend to 3D applications, as token sequence length would grow cubically, exceeding our computational budget. Potential remedies for 3D scaling include: (1) using VAE-like techniques/transformers to compress and evolve the fields in the latent space, then applying attention mechanisms to decode at any location [50, 14], and (2) employing crop-like methods to learn evolution of fixed-width windows similar to CNNs. These approaches could potentially be combined to address the cubic scaling challenge.
Since VICON is an autoregressive model, it still suffers from error accumulation of the final pair’s condition frame during rollout. Future work could explore resolving error accumulation through techniques such as injecting Gaussian noise or using push-forward methods [51].
The channel-union approach is also limited when incorporating domains beyond fluid dynamics with fundamentally different state variables. Recent advances in inter-channel attention mechanisms [52] offer promising directions for handling channel scaling issues when expanding to more diverse physics systems.
Finally, adapting VICON to handle irregular domains such as graphs or meshes would broaden its applications to areas like solid mechanics and molecular dynamics. The original ICON framework has demonstrated success across various problem types including inverse problems and steady-state tasks [25], suggesting potential for extending VICON beyond forward temporal prediction. Addressing these challenges will advance the promising paradigm of in-context learning for a broader range of physical systems.
Acknowledgments
@@ A.1 PDEArena-Incomp Dataset
The space-time domain is $0,32]^{2}\times[18,105]$ where $dt=1.5$ and the space resolution is $128\times 128$. A scalar particle field is being transported with the fluids. The velocity fields satisfy Dirichlet boundary condition, and the scalar field satisfies Neumann boundary condition. The forcing term $\bm{f}$ is randomly sampled. The quantities of interest are the velocities and the scalar particle field.
@@ B.1 VICON Model Details
Here we provide key architectural parameters of VICON model implementation in table˜4.
@@ B.2 Training Details
We implement our method in PyTorch [53] and utilize data parallel training [54] across two NVIDIA RTX 4090 GPUs. We employ the AdamW optimizer with a cosine learning rate schedule that includes a linear warmup phase. We apply gradient clipping with a maximum norm of 1.0 to ensure training stability. All optimization parameters are detailed in table˜5.
@@ 2 S $\leftarrow$ [ ]
We demonstrate two rollout strategies in Tables 6 and 7, showing single-step and flexible-step strategies, respectively. In both cases, the initial frames span from time step 0 to 9, and we aim to predict the trajectory up to time step 20.
@@
As shown in table˜7, the flexible-step strategy initially uses smaller strides to build sufficient examples, then employs maximum strides for later rollouts, as detailed in section˜4.5. We note that repeated in-context examples appear in table˜7, which is common when the maximum stride is large and the initial frames cannot form enough examples. While the number of in-context examples in VICON is flexible and our model can accommodate fewer examples than the designed length, our preliminary experiments indicate that the model performs better with more in-context examples, even when some examples are repeated.
@@ 30   Pdt $\leftarrow$ P[s]
Tables 8 and 9 illustrate our adaptive strategies when frames 2, 5, and 9 are missing from the initial sequence (with all other settings identical to those in section˜C.1). The tables demonstrate single-step and flexible-step (max stride: 3) approaches, respectively, highlighting how our algorithm dynamically selects appropriate example pairs to maintain prediction capability despite missing some temporal samplings.
@@ D.1 Ablation Studies
We conducted ablation studies on patch resolution by varying patch sizes (8, 16, 32, 64) to find a balance between spatial granularity and computational resource constraint. While smaller patches theoretically capture finer details, they generate longer token sequences, hitting memory caps due to transformer’s quadratic complexity. For patches smaller than $16\times 16$, we had to reduce token dimensions (1024→512) and feedforward dimensions (from 2048 to 1024), while patches below $8\times 8$ remained infeasible even with mixed precision training. fig.˜4(a-c) shows that patch size 16 achieves optimal performance across all datasets—balancing sequence length and representational capacity. Performance degradation with coarser patches (32, 64) aligns with expectations, while degradation with finer patches stems from the necessary reduction in hidden dimensions, highlighting a fundamental challenge in scaling to 3D applications.
@@
We evaluated different combinations of positional encodings as shown in fig.˜4(d-f). Our architecture employs two types of encodings: patch position encodings (for spatial relationships between patches) and function encodings (differentiating input/output in different pairs) (section˜4.2).
@@
We investigated how the number of in-context examples affects model performance, as shown in fig.˜4(g-i). Following insights from the original ICON work [25], we evaluated context lengths of 6, 8, and 10 pairs, balancing performance against computational efficiency. The results show that 10 in-context examples deliver significantly better results for PDEBench-Comp-LowVis compared to shorter contexts, while for the remaining datasets, 10 and 8 in-context examples perform similarly (both outperforming 6 examples).
D.2 Analysis of Checkerboard Artifacts in PDEBench-Comp-HighVis
In our experiments, we observed block-like artifacts in the error/difference maps for the PDEBench-Comp-HighVis dataset that align with the model’s 16×16 patch structure. Through detailed analysis, we identified two contributing factors: 1) drastic field value variation across the trajectory time horizon, and 2) a training-inference normalization mismatch.
Drastic field value variation across trajectory time horizon. High-viscosity flows rapidly dissipate kinetic energy, causing velocity fields to become nearly stationary after approximately 10 frames. Our analysis of standard deviations across different time ranges reveals these differences:
Table 10: Standard deviation analysis across time phases for PDEBench datasets. The bolded velocity channels show PDEBench-Comp-HighVis has extreme variation between early and late phases, unlike PDEBench-Comp-LowVis with nearly consistent variation. This dynamic difference makes training a uniformed model particularly challenging.
Dataset Velocity u Velocity v Pressure
All 1-10 >10 All 1-10 >10 All
PDEBench-Comp-HighVis 0.359 0.511 0.091 0.351 0.506 0.049 14.254
PDEBench-Comp-LowVis 0.989 1.104 0.873 1.036 1.165 0.903 29.687
Training-inference normalization mismatch. During training, context pairs are sampled uniformly across all frames. During inference, context pairs are sampled from initial frames only. As shown in table˜10, the standard deviation at these two different time ranges are vastly different which results in normalization mismatch and exacerbates prediction errors in the already-challenging problem. Future work could explore using global statistics from multiple trajectories instead of single-trajectory instance normalization during deployment. Other potential remedies include using overlapping patches or moving patch positions during rollout, though these represent engineering improvements orthogonal to our core contributions.
D.3 More Results and Visualizations
@@
Step 1
[1e-2]
PDEArena-Incomp 11.01 11.01 5.97 6.17
@@
Step 5
[1e-2]
PDEArena-Incomp 22.68 20.62 11.70 14.81
@@
Step 10
[1e-2]
PDEArena-Incomp 35.67 30.53 20.74 27.89
@@
Last step
[1e-2]
PDEArena-Incomp 76.77 68.03 65.27 93.52
@@
All average
[1e-2]
PDEArena-Incomp 56.27 48.50 41.20 55.95
@@
Step 1
[1e-2]
PDEArena-Incomp 5.63 5.63 3.02 3.12
@@
Step 5
[1e-2]
PDEArena-Incomp 10.38 9.46 5.37 6.79
@@
Step 10
[1e-2]
PDEArena-Incomp 14.65 12.55 8.55 11.49
@@
Last Step
[1e-2]
PDEArena-Incomp 16.50 14.56 14.03 19.47
@@
All average
[1e-2]
PDEArena-Incomp 16.26 14.31 11.86 15.77
@@
Tables 11 and 12 summarize the relative and absolute $L^{2}$ rollout errors across different timesteps for all evaluated models on the three datasets.
@@
fig.˜5 illustrates VICON’s performance with varying timestep strides ($s_{\text{max}}=1,3,5,7$) across all three datasets, demonstrating the model’s ability to adapt to different temporal resolutions without retraining.
@@
fig.˜6 displays the visualization of TKE fields for ground truth and model predictions, highlighting VICON’s superior ability to preserve critical turbulent flow structures compared to baseline approaches.
@@
Step 1
PDEArena-Incomp 11.10 9.66
@@
Step 5
PDEArena-Incomp 23.00 24.09
@@
Step 10
PDEArena-Incomp 36.18 41.66
@@
Last Step
PDEArena-Incomp 77.81 123.0
@@
All average
PDEArena-Incomp 56.26 76.35
@@
fig.˜7 and table˜13 quantify the performance differences between jointly trained and separately trained models, demonstrating the significant advantages of multi-physics training across all datasets.
@@
table˜14 summarizes the computational resources required by each method, showing VICON’s advantages in terms of training cost, inference speed, and model parameter count compared to both DPOT and MPP baselines.
@@
Figure 8: Comparison of rollout errors (scale by std) with different input data noise levels across all datasets. Each column (left to right) shows VICON, DPOT, and MPP models, while each row (top to bottom) represents PDEBench-Comp-LowVis, PDEBench-Comp-HighVis, and PDEArena-Incomp datasets. For MPP and DPOT which require fixed dtdt and context window, interpolation is used to generate missing frames, while VICON can directly handle irregular temporal data without interpolation.
@@
fig.˜8 show the rollout results with imperfect temporal measurements.
@@
We compare the output of different models in fig.˜9, fig.˜10, and fig.˜11. Figures 12 and 13 present additional visualizations of the VICON model outputs compared to ground truth and baseline predictions, highlighting the qualitative advantages of our approach.